Two copper balls each weighining 10g are kept in air 1 cm apart If one electron from every 10^6 atoms is transfered from one ball to the other the coloumb force between them is

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

molecular weight = 63.55

so mass of 1 mole of cu is =63.55

 in 63.55 g number of cu atom   = 6.022 X 1023 

 in 10 g number of cu atom   = 10*6.022/63.55  X 1023 =1.00 X 10 23

  number of electron tranfered  = 10 23   / 106

                                             =1017

so charged genereted in one ball is q  =1.6 X 10-19  X 1017

                                                   =1.6 X 10-2 C

so force F = 1/4∏ε qq/r2

              =9 X 109 * (1.6 X 10-2 )2 / 10-4

              =2.304 1010 N

 Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem below.
 
molecular weight = 63.55
so mass of 1 mole of cu is =63.55
in 63.55 g number of cu atom   = 6.022 X 1023  
in 10 g number of cu atom   = 10*6.022/63.55  X 1023 =1.00 X 10 23
  number of electron tranfered  = 10 23   / 106  = 1017
so charged genereted in one ball is q  =1.6 X 10-19  X 1017=1.6 X 10-2 C
so force F = 1/4∏ε qq/r2 =9 X 109 * (1.6 X 10-2 )2 / 10-4
                                    =2.304 1010 N
 
Thanks and regards,
Kushagra